walid is four times as old as sami, five years ago,walid was 7 times old as sami. how old is walid snd how old is sami ?
Answer by stanbon(75887)

walid is four times as old as sami, five years ago,walid was 7 times old as sami. how old is walid snd how old is sami ?
Equations:
w = 4s
w-5 = 7(s-5)
----
Substitute for "w" and solve for "s"::
4s-5 = 7s-35
3s = 30
s = 10 (sami's age now)
w = 4s = 40 (walid's age now)
--------------
Cheers,
Stan H.
--------------